OpenStudy (anonymous):

Find the area of a right triangle with a base length of 3 units, and a hypotenuse of 5.

OpenStudy (anonymous):

OKay so the first step is to find what the other leg is (a). \[a^2+3^2=5^2\] \[a^2+9=25\] a=16.

OpenStudy (anonymous):



OpenStudy (anonymous):

so what is the area of the triangle?

OpenStudy (anonymous):

I am trying to find the h then * by 2

OpenStudy (anonymous):